Let {xn} n=1to ∞ and {yn}n=1to ∞ be sequences of positive numbers such that limn→∞ xn/yn = 0. Then show that if {yn}n=1to ∞ is bounded, then limn→∞ xn = 0



1
Expert's answer
2021-02-24T06:08:01-0500

We have given that "{x_n}" n=1 to ∞ and "{y_n}\ufeff" n=1to ∞ be sequences of positive numbers such that

"lim_{n\rightarrow\infty} \dfrac{x_n}{y_n}"

It is also given that "y_n" is a bounded sequence

Hence "y_n" is not equal to zero

therefore in the expression "lim_{n\rightarrow \infty }\dfrac{x_n}{y_n}"

"y_n \neq 0"

Therefore,

"lim_{n\rightarrow \infty}x_n = 0"